# Soft Prompt Generation for Domain Generalization

### v1

@@ Abstract
Large pre-trained vision language models (VLMs) have shown impressive zero-shot ability on downstream tasks with manually designed prompt, which are not optimal for specific domains. To further adapt VLMs to downstream tasks, soft prompt is proposed to replace manually designed prompt, which acts as a learning vector that undergoes fine-tuning based on specific domain data. Prior prompt learning methods primarily learn a fixed prompt and residuled prompt from training samples. However, the learned prompts lack diversity and ignore information about unseen domains, potentially compromising the transferability of the prompts. In this paper, we reframe the prompt learning framework from a generative perspective and propose a simple yet efficient method for the Domain Generalization (DG) task, namely Soft Prompt Generation (SPG). To the best of our knowledge, we are the first to introduce the generative model into prompt learning in VLMs and explore its potential for producing soft prompts by relying solely on the generative model, ensuring the diversity of prompts. Specifically, SPG consists of a two-stage training phase and an inference phase. During the training phase, we introduce soft prompt labels for each domain, aiming to incorporate the generative model domain knowledge. During the inference phase, the generator of the generative model is employed to obtain instance-specific soft prompts for the unseen target domain. Extensive experiments on five domain generalization benchmarks of three DG tasks demonstrate that our proposed SPG achieves state-of-the-art performance. The code will be available soon.
@@
Equal contribution.
Corresponding author.
@@ 1 Introduction
Large vision language models, such as CLIP [28] and ALIGN [14], have attracted significant attention owing to their effective adaptation to downstream tasks with manually designed prompts. However, manually designed prompts are not always optimal for domain-specific tasks. Instead of manually designed prompts, the soft prompt [35] can be optimized in a data-driven manner through backpropagation. The soft prompt, serving as a learning vector, is refined through fine-tuning with domain-specific data to better adapt to downstream tasks, including base-to-novel generalization [35, 16] and domain adaptation [11, 2]. Despite the progress made in prompt learning, generalization performance tends to decline significantly when facing distribution shifts.
Many efforts have been made to address the distribution shift problem by domain generalization (DG) [33, 22]. The objective of DG is to train a model using data from one or multiple related but distinct source domains in a manner that enables the model to generalize effectively to any out-of-distribution (OOD) unseen target domain. Due to the extensive learning in various domains, vision language models like CLIP exhibit a generalization performance that can even rival state-of-the-art traditional supervised learning algorithms. Taking a step, many works adopt prompt learning techniques [26, 34, 31] to enhance the generalization performance of CLIP. Niu et al. [26] introduce domain bank to incorporate textual domain knowledge into soft prompts. Both Zhou et al. [34] and Zhang et al. [31] establish a lightweight neural network (i.e., multilayer perceptron) to generate soft prompts conditioned on images with a residual or concatenation manner. However, these methods either lack diversity and visibility into the information from the target domain or rely on simple models to approximate the relationship between images and soft prompts. Consequently, the learned prompts may still fall short of being transferable.
@@ 4 Method
The generator of CGAN is employed to produce domain-specific soft prompts for each image of the target domain. The probability of the image belonging to the $i$th class can be formulated as:
@@ 5.1 Experimental Setting
Experiments are conducted on five popular benchmark datasets of domain generalization, namely PACS [18], VLCS [9], OfficeHome [29], TerraIncognita (TerraInc.) [3] and DomainNet [27].
@@ 5.2 Comparison with SOTA Methods
Table 2 shows the experimental results of the leave-all-but-one-domain-out performance of PACS, VLCS, Office-Home, TerraIncognita and DomainNet for ResNet50 backbone. SPG outperforms other CLIP-based fine-tuning methods on five datasets. We observe averaged improvements of 19.0% and 1.4% compared with linear probing of CLIP and the SOTA method CoOp, respectively. This demonstrates that by leveraging the generative model for prompt generation, SPG is able to produce more domain-relevant and adaptive prompts, leading to improved performance across different domains.
@@ 0.A.0.5 DomainNet [27]
Appendix 0.B Implementation Details
@@
Appendix 0.C Supplement Experiments
The supplement experiments mainly demonstrate the domain-wise results of single-source DG and multi-source DG.
0.C.1 Multi-source DG Comparisons
@@
0.C.2 Single-source DG Comparisons

### v2

@@ Abstract
Large pre-trained vision language models (VLMs) have shown impressive zero-shot ability on downstream tasks with manually designed prompt. To further adapt VLMs to downstream tasks, soft prompt is proposed to replace manually designed prompt, which undergoes fine-tuning based on specific domain data. Prior prompt learning methods primarily learn a fixed prompt or residuled prompt from training samples. However, the learned prompts lack diversity and ignore information about unseen domains. In this paper, we reframe the prompt learning framework from a generative perspective and propose a simple yet efficient method for the Domain Generalization (DG) task, namely Soft Prompt Generation (SPG). Specifically, SPG consists of a two-stage training phase and an inference phase. During the training phase, we introduce soft prompt label for each domain, aiming to incorporate the generative model domain knowledge. During the inference phase, the generator of the generative model is employed to obtain instance-specific soft prompts for the unseen target domain. Extensive experiments on five domain generalization benchmarks of three DG tasks demonstrate that SPG achieves state-of-the-art performance. The code is available at https://github.com/renytek13/Soft-Prompt-Generation-with-CGAN.
@@
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Large vision language models, such as CLIP [32] and ALIGN [16], have attracted significant attention owing to their effective adaptation to downstream tasks with manually designed prompts. However, manually designed prompts are not always optimal for domain-specific tasks. Instead of manually designed prompts, the soft prompt [40, 31, 19] can be optimized in a data-driven manner through backpropagation. The soft prompt, serving as a learning vector, is refined through fine-tuning with domain-specific data to better adapt to downstream tasks, including base-to-novel generalization [40, 18] and domain adaptation [13, 2, 35]. Despite the progress made in prompt learning, generalization performance tends to decline significantly when facing distribution shifts.
Many efforts have been made to address the distribution shift problem by domain generalization (DG) [38, 25]. The objective of DG is to train a model using data from one or multiple related but distinct source domains in a manner that enables the model to generalize effectively to any out-of-distribution (OOD) unseen target domain. Due to the extensive learning in various domains, vision language models like CLIP exhibit a generalization performance that can even rival state-of-the-art traditional supervised learning algorithms. Taking a step, many works adopt prompt learning techniques [29, 39, 36, 7, 6] to enhance the generalization performance of CLIP. Niu et al. [29] introduce domain bank to incorporate textual domain knowledge into soft prompts. Both Zhou et al. [39] and Zhang et al. [36] establish a lightweight neural network (i.e., multilayer perceptron) to generate soft prompts conditioned on images with a residual or concatenation manner. However, these methods either lack diversity and visibility into the information from the target domain or rely on simple models to approximate the relationship between images and soft prompts. Consequently, the learned prompts may still fall short of being transferable.
@@ 4 Method
The generator of CGAN is employed to produce a domain-specific soft prompt for each image of the target domain. The probability of the image belonging to the $i$th class can be formulated as:
@@ 5.1 Experimental Setting
Experiments are conducted on five popular benchmark datasets of DG, namely PACS [21], VLCS [11], OfficeHome [33], TerraIncognita (TerraInc.) [3] and DomainNet [30]. The dataset details can be seen in the appendix.
@@ 5.2 Comparison with SOTA Methods
Table 2 shows the experimental results of the leave-all-but-one-domain-out performance of PACS, VLCS, Office-Home, TerraIncognita and DomainNet for ResNet50 backbone. SPG outperforms other CLIP-based fine-tuning methods on five datasets. We observe averaged improvements of 19.0% and 1.4% compared with linear probing of CLIP and the SOTA method CoOp, respectively. This demonstrates that by leveraging the generative model for prompt generation, SPG is able to produce more domain-relevant and adaptive prompts, leading to improved performance across different domains. The domain-wise results for both multi-source DG and single-source DG are provided in the appendix.
@@ 0.A.0.5 DomainNet [30]
Appendix 0.B Algorithm Details
The overall framework of the pseudo-code of SPG is described in Algorithm 1 and Algorithm 2. Algorithm 1 demonstrates the process of stage I: Domain Prompt Labels Learning, and Algorithm 2 shows the process of stage II: Generative Model Pre-training.
Algorithm 1 Soft Prompt Generation - Domain Prompt Labels Learning
Requirement: pre-defined ${N}_{c}$ class names in the target task
Input: images and labels of training samples $(\mathbf{x}^{{d}_{i}}_{j},{y}^{{d}_{i}}_{j})$, number of training iterations $L$
Output: ${N}_{d}$ domain prompt labels
1:# learn prompt label on each domain separately
2:for $i=1,2,\ldots,{N}_{d}$ do
3:      # initialize $i$-th domain prompt label with prompt prefix $\mathbf{v}^{p}$ and learnable vector $\mathbf{v}^{i}$.
4:     $\mathbf{v}^{{d}_{i}}\leftarrow\mathtt{initialize}(\mathbf{v}^{p},\mathbf{v}^{i})$
5:      # $L$ training iterations for learning each domain prompt label
6:     for $\mathrm{iteration}=1,2,\ldots,L$ do
7:            # update learnable vector ${v}^{d_{i}}$
8:         ${\mathbf{v}^{d_{i}}}^{*}=\arg\min_{\mathbf{v}}\mathbb{E}_{\mathbf{x}^{d_{i}}_{j},{y}^{d_{i}}_{j}}[-\log p({y}_{j}^{d_{i}}\mid\mathbf{x}^{d_{i}}_{j},\mathbf{v}^{d_{i}})]$
9:         Update $\mathbf{v}^{d_{i}}$ by gradient descent
10:     end for
11:end for
12:Store optimal domain prompt label $\mathbf{v}^{d_{i}}$ for each domain
Algorithm 2 Soft Prompt Generation - Generative Model Pre-training
Requirement: A CGAN with a generator $G$ and a discriminator $D$, real vector $\mathbf{v}_{\mathrm{real}}$ and fake vector $\mathbf{v}_{\mathrm{fake}}$, and domain prompt labels $\mathbf{v}^{d_{i}}$
Input: image embeddings $f(\mathbf{x})$, number of training iterations $L$
Output: optimal prompt for each image
1:# $L$ training iterations
2:for $\mathrm{iteration}=1,2,\ldots,L$ do
3:       # define input noise variable $\mathbf{z}$ and combines these noise variables with image embeddings $f(\mathbf{x})$ as input of generator G and output $\mathbf{v}_{g}$
4:     $\mathbf{v}_{g}=G(input)\leftarrow input=\mathtt{concat}([\mathbf{z},f(\mathbf{x})])$
5:       # discriminator determines the authenticity of domain prompt labels $\mathbf{v}^{d_{i}}$ and generated prompt $\mathbf{v}_{g}$
6:     $\mathbf{v}_{d,\mathrm{real}}=D(\mathbf{v}^{d_{i}})$, $\mathbf{v}_{d,\mathrm{fake}}=D(\mathbf{v}_{g})$
7:       # compute $\mathcal{L}_{\mathrm{real}}$ with discriminator_real output $\mathbf{v}_{d,\mathrm{real}}$ and pre-defined real vector $\mathbf{v}_{\mathrm{real}}$
8:     $\mathcal{L}_{\mathrm{real}}\leftarrow\mathtt{mse\_loss}(\mathbf{v}_{d,\mathrm{real}},\mathbf{v}_{\mathrm{real}})$
9:       # compute $\mathcal{L}_{\mathrm{fake}}$ with discriminator_fake output $\mathbf{v}_{d,\mathrm{fake}}$ and pre-defined fake vector $\mathbf{v}_{\mathrm{fake}}$
10:     $\mathcal{L}_{\mathrm{fake}}\leftarrow\mathtt{mse\_loss}(\mathbf{v}_{d,\mathrm{fake}},\mathbf{v}_{\mathrm{fake}})$
11:     $\mathcal{L}_{\mathrm{discriminator}}\leftarrow\mathcal{L}_{\mathrm{real}}+\mathcal{L}_{\mathrm{fake}}$
12:     Update parameters of $D$ using $\mathcal{L}_{\mathrm{discriminator}}$ by gradient descent
13:       # compute $\mathcal{L}_{\mathrm{generator}}$ with discriminator_fake output $\mathbf{v}_{d,\mathrm{fake}}$ and pre-defined real vector $\mathbf{v}_{\mathrm{real}}$
14:     $\mathcal{L}_{\mathrm{generator}}\leftarrow\mathtt{mse\_loss}(\mathbf{v}_{d,\mathrm{fake}},\mathbf{v}_{\mathrm{real}})$
15:     Update parameters of $G$ using $\mathcal{L}_{\mathrm{generator}}$ by gradient descent
16:end for
17:Generate the optimal prompt for each input image
Appendix 0.C Implementation Details
@@
Appendix 0.D Supplement Experiments
The supplement experiments mainly demonstrate the domain-wise results of multi-source DG and single-source DG.
0.D.1 Multi-source DG Comparisons
@@
0.D.2 Single-source DG Comparisons
@@
0.D.3 Ablation Study
Component ablation. The domain prompt label and generative model are indispensable components of our SPG method and cannot be directly removed, but can be replaced. For domain prompt label, We replace the text prompt used in our work with VP [1]. For backbone, we replace the CGAN [27] with an MLP. Additional domain label and backbone ablations are as follows.
Table 17: Component ablation for multi-source DG in terms of mean leave-one-domain-out performance with ResNet50 as the backbone. Bold denotes the best scores.
Examples PACS VLCS O.H. TerraInc. Do.Net Avg
w/ VP 89.4 79.8 67.8 19.1 44.7 60.2
w/ MLP 90.4 79.6 72.3 31.7 47.8 64.4
SPG (Ours) 92.8 84.0 73.8 37.5 50.1 67.6